# On the GKZ discriminant locus

### v3

@@ Abstract.
The first author is supported by a MIS grant from the National Fund for Scientific Research (FNRS) and an ACR grant from the Université Libre de Bruxelles. The second author is a senior researcher at the Research Foundation Flanders (FWO). While working on this project he was supported by the ERC grant SCHEMES and FWO grant G0D8616N: “Hochschild cohomology and deformation theory of triangulated categories”.
@@ 1. Introduction and statement of the main result
Let $k,d\in{\mathbb{N}}$ and let $A=\{a_{1},\dots,a_{d}\}\subset{\mathbb{Z}}^{k}$. Let $P\subset{\mathbb{R}}^{k}$ be the convex hull of $(a_{i})_{i}$. Let $F$ be a face of $P$. Denote by $\nabla_{F}\subset{\mathbb{C}}^{F\cap A}$ the discriminant locus [^1] In [GKZ08, §9.1.A] $\nabla_{F}$ is defined as the closure of our $\nabla_{F}$. of $f^{F}:=\sum_{a_{i}\in F}\alpha_{i}x^{a_{i}}\in{\mathbb{C}}[x_{1}^{\pm 1},\ldots,x_{k}^{\pm 1}]$, i.e. the set of all $\alpha\in{\mathbb{C}}^{F\cap A}$ such that the singular locus of $V(f_{\alpha}^{F})$ (i.e. the common zeros of $f_{\alpha}^{F}$, $(\partial_{i}f_{\alpha}^{F})_{i=1,\ldots,k}$) meets (*$({\mathbb{C}}^{*})^{k}$ [GKZ08, §9.1.A]. It is known that ¯$\bar{\nabla}_{F}$ is irreducible and $\dim\bar{\nabla}_{F}<|F\cap A|$ [GKZ08, Chapter 9, after (1.1)]. Set $V(A):=\bigcup_{F}p^{-1}_{F}(\nabla_{F})$, where $p_{F}:{\mathbb{C}}^{A}\to{\mathbb{C}}^{F\cap A}$ is the projection. If ¯$\bar{\nabla}_{F}$ is a hypersurface in ${\mathbb{C}}^{F\cap A}$ (i.e. if $\dim\nabla_{F}=|F\cap A|-1$) then we denote by $\Delta_{F}$ its defining equation, otherwise we set $\Delta_{F}=1$. We denote the composition $\Delta_{F}\circ p_{F}:{\mathbb{C}}^{A}\rightarrow{\mathbb{C}}$ also by $\Delta_{F}$. The principal $A$-determinant is a certain polynomial function on ${\mathbb{C}}^{A}$ which is defined in [GKZ08, §10.1.A]. In [GKZ08, Theorem 10.1.2] it is shown that $E_{A}=\prod_{F\subseteq P}\Delta_{F}^{m_{F}}$ for suitable multiplicities $m_{F}\geq 1$. Hence
@@ 2. Toric preliminaries
[GKZ08, Proposition 5.1.9] The torus orbits in $X_{A}$ are in $1{:}1$ correspondence with the faces of $P$. The orbit $O_{F}$ corresponding to a face $F$ of $P$ consists of $\{(z_{1}:\dots:z_{d})\in X_{A}\mid z_{i}\neq 0\;\text{if $a_{i}\in F$},z_{i}=0\;\text{if $a_{i}\not\in F$}\}$. $O_{F}$ is also equal to $\{(z_{1}:\dots:z_{d})\in{\mathbb{P}}^{d-1}\mid\exists u\in({\mathbb{C}}^{*})^{k}\text{such that}\;z_{i}=u^{a_{i}}\;\text{if $a_{i}\in F$}\text{ and }z_{i}=0\;\text{if $a_{i}\not\in F$}\}$.
@@
The first two claims are proved in loc. cit. As the last claim does not seem to appear explicitly in the given reference, we provide a proof here, assuming the other claims. We first claim that the point $z_{c}:=(z_{1}:\dots:z_{d})$ with $z_{i}=1$ if $a_{i}\in F$ and $z_{i}=0$ if $a_{i}\not\in F$ is contained in $O_{F}$. We first prove $z_{c}\in X_{A}$. Let $\lambda\in{\mathbb{Z}}^{k}$, $c\in{\mathbb{Z}}$ be such that $\langle\lambda,a_{i}\rangle>c$ if $a_{i}\not\in F$ and $\langle\lambda,a_{i}\rangle=c$ if $a_{i}\in F$ (such $\lambda$ exists because $F$ is a face of the convex hull of $A$). Let $(y_{t})_{t}$ be the $1$-parameter subgroup of (*$({\mathbb{C}}^{*})^{k}$ defined by $\lambda$, i.e. $y_{t}=(t^{\lambda_{1}},\dots,t^{\lambda_{k}})$. Put $z_{0}:=(1:\dots:1)\in X_{A}$. Then $\lim_{t\to 0}y_{t}\cdot z_{0}=\lim_{t\to 0}t^{-c}y_{t}\cdot z_{0}=\lim_{t\to 0}(t^{\langle\lambda,a_{1}\rangle-c}:\dots:t^{\langle\lambda,a_{d}\rangle-c})=z_{c}$ implies $z_{c}\in X_{A}$. Now the description of $O_{F}$ given in the statement of the proposition implies $z_{c}\in O_{F}$.
@@ 3.1. Discriminant locus and finiteness
On the other hand, if $\alpha\in V(A)=\bigcup_{F}p_{F}^{-1}(\nabla_{F})$ then there exists a face $F$ of $P$, $u\in({\mathbb{C}}^{*})^{k}$ such that $(\partial_{i}f_{\alpha}^{F})(u)=0$ for $i=1,\ldots,k$. By Proposition 2.1 the point $w$ with homogeneous coordinates $w_{i}=u^{a_{i}}$ if $a_{i}\in F$ and $w_{i}=0$ otherwise is contained in $X_{A}$. The point with the same (affine) coordinates in the cone over $X_{A}$ corresponds by Proposition 2.2 to a ring homomorphism $\delta:R\to{\mathbb{C}}$ which sends $x^{a_{i}}$ to $w_{i}$ for $i=1,\ldots,d$, i.e. $x^{a_{i}}$ to $u^{a_{i}}$ if $a_{i}\in F$ and to 00 otherwise. It follows that $(\delta\gamma)(y_{i})=0$, $1\leq i\leq k$. On the other hand $(w_{i})_{i}\neq(0,\ldots,0)$ since $(w_{i})_{i}$ corresponds to a point in ${\mathbb{P}}^{d-1}$. Hence $\operatorname{Spec}\bar{R}$ has a closed point $\operatorname{ker}\delta$ different from $\bar{m}$ and hence $\bar{R}$ must be infinite dimensional. ∎

### v4

@@ Abstract.
The first author is supported by a MIS grant from the National Fund for Scientific Research (FNRS) and an ACR grant from the Université Libre de Bruxelles. The second author is a senior researcher at the Research Foundation Flanders (FWO). This project has received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (grant agreement No 885203). The project was also supported by FWO grant G0D8616N: “Hochschild cohomology and deformation theory of triangulated categories”.
@@ 1. Introduction and statement of the main result
Let $k,d\in{\mathbb{N}}$ and let $A=\{a_{1},\dots,a_{d}\}\subset{\mathbb{Z}}^{k}$. Let $P\subset{\mathbb{R}}^{k}$ be the convex hull of $(a_{i})_{i}$. Let $F$ be a face of $P$. Denote by $\nabla_{F}\subset{\mathbb{C}}^{F\cap A}$ the discriminant locus [^1] In [GKZ08, §9.1.A] $\nabla_{F}$ is defined as the closure of our $\nabla_{F}$. of $f^{F}:=\sum_{a_{i}\in F}\alpha_{i}x^{a_{i}}\in{\mathbb{C}}[x_{1}^{\pm 1},\ldots,x_{k}^{\pm 1}]$, i.e. the set of all $\alpha\in{\mathbb{C}}^{F\cap A}$ such that the singular locus of $V(f_{\alpha}^{F})$ (i.e. the common zeros of $f_{\alpha}^{F}$, $(\partial_{i}f_{\alpha}^{F})_{i=1,\ldots,k}$) meets $({\mathbb{C}}^{*})^{k}$ [GKZ08, §9.1.A]. It is known that $\bar{\nabla}_{F}$ is irreducible and $\dim\bar{\nabla}_{F}<|F\cap A|$ [GKZ08, Chapter 9, after (1.1)]. Set $V(A):=\bigcup_{F}p^{-1}_{F}(\nabla_{F})$, where $p_{F}:{\mathbb{C}}^{A}\to{\mathbb{C}}^{F\cap A}$ is the projection. If $\bar{\nabla}_{F}$ is a hypersurface in ${\mathbb{C}}^{F\cap A}$ (i.e. if $\dim\nabla_{F}=|F\cap A|-1$) then we denote by $\Delta_{F}$ its defining equation, otherwise we set $\Delta_{F}=1$. We denote the composition $\Delta_{F}\circ p_{F}:{\mathbb{C}}^{A}\rightarrow{\mathbb{C}}$ also by $\Delta_{F}$. The principal $A$-determinant is a certain polynomial function on ${\mathbb{C}}^{A}$ which is defined in [GKZ08, §10.1.A]. In [GKZ08, Theorem 10.1.2] it is shown that $E_{A}=\prod_{F\subseteq P}\Delta_{F}^{m_{F}}$ for suitable multiplicities $m_{F}\geq 1$. Hence
@@ 2. Toric preliminaries
[GKZ08, Proposition 5.1.9] The torus orbits in $X_{A}$ are in 1:11{:}1 correspondence with the faces of $P$. The orbit $O_{F}$ corresponding to a face $F$ of $P$ consists of $\{(z_{1}:\dots:z_{d})\in X_{A}\mid z_{i}\neq 0\;\text{if $a_{i}\in F$},z_{i}=0\;\text{if $a_{i}\not\in F$}\}$. $O_{F}$ is also equal to $\{(z_{1}:\dots:z_{d})\in{\mathbb{P}}^{d-1}\mid\exists u\in({\mathbb{C}}^{*})^{k}\text{such that}\;z_{i}=u^{a_{i}}\;\text{if $a_{i}\in F$}\text{ and }z_{i}=0\;\text{if $a_{i}\not\in F$}\}$.
@@
The first two claims are proved in loc. cit. As the last claim does not seem to appear explicitly in the given reference, we provide a proof here, assuming the other claims. We first claim that the point $z_{c}:=(z_{1}:\dots:z_{d})$ with $z_{i}=1$ if $a_{i}\in F$ and $z_{i}=0$ if $a_{i}\not\in F$ is contained in $O_{F}$. We first prove $z_{c}\in X_{A}$. Let $\lambda\in{\mathbb{Z}}^{k}$, $c\in{\mathbb{Z}}$ be such that $\langle\lambda,a_{i}\rangle>c$ if $a_{i}\not\in F$ and $\langle\lambda,a_{i}\rangle=c$ if $a_{i}\in F$ (such $\lambda$ exists because $F$ is a face of the convex hull of $A$). Let $(y_{t})_{t}$ be the $1$-parameter subgroup of $({\mathbb{C}}^{*})^{k}$ defined by $\lambda$, i.e. $y_{t}=(t^{\lambda_{1}},\dots,t^{\lambda_{k}})$. Put $z_{0}:=(1:\dots:1)\in X_{A}$. Then $\lim_{t\to 0}y_{t}\cdot z_{0}=\lim_{t\to 0}t^{-c}y_{t}\cdot z_{0}=\lim_{t\to 0}(t^{\langle\lambda,a_{1}\rangle-c}:\dots:t^{\langle\lambda,a_{d}\rangle-c})=z_{c}$ implies $z_{c}\in X_{A}$. Now the description of $O_{F}$ given in the statement of the proposition implies $z_{c}\in O_{F}$.
@@ 3.1. Discriminant locus and finiteness
On the other hand, if $\alpha\in V(A)=\bigcup_{F}p_{F}^{-1}(\nabla_{F})$ then there exists a face $F$ of $P$, $u\in({\mathbb{C}}^{*})^{k}$ such that $(\partial_{i}f_{\alpha}^{F})(u)=0$ for $i=1,\ldots,k$. By Proposition 2.1 the point $w$ with homogeneous coordinates $w_{i}=u^{a_{i}}$ if $a_{i}\in F$ and $w_{i}=0$ otherwise is contained in $X_{A}$. The point with the same (affine) coordinates in the cone over $X_{A}$ corresponds by Proposition 2.2 to a ring homomorphism $\delta:R\to{\mathbb{C}}$ which sends $x^{a_{i}}$ to $w_{i}$ for $i=1,\ldots,d$, i.e. $x^{a_{i}}$ to $u^{a_{i}}$ if $a_{i}\in F$ and to 0 otherwise. It follows that $(\delta\gamma)(y_{i})=0$, $1\leq i\leq k$. On the other hand $(w_{i})_{i}\neq(0,\ldots,0)$ since $(w_{i})_{i}$ corresponds to a point in ${\mathbb{P}}^{d-1}$. Hence $\operatorname{Spec}\bar{R}$ has a closed point $\operatorname{ker}\delta$ different from $\bar{m}$ and hence $\bar{R}$ must be infinite dimensional. ∎